Circular No. 8992

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE REDDENED SUPERNOVAE IN NGC 972 AND NGC 5900

C. C. Dudley and J. Fischer, Naval Research Laboratory, report the discovery of two possible high-extinction supernovae in images obtained with the U.S. Naval Observatory 1.55-m telescope (+ 2.1-μm K′ filter + ASTROCAM 1024×1024 ALADDIN InSb imager) at Flagstaff (discovery observations tabulated below, with discovery dates and approximate K′ magnitudes at time of discovery).

PSN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
K0210-1 2002 Oct. 16.3 2^{h}34^{m}13.^{s}34 +29^{o}18′36′′.2 14.7 5′′ S
K0304-1 2003 Apr. 17.4 15 15 06.97 +42 12 19.3 17.7 20′′ E, 17′′ S

Additional approximate K′ magnitudes for PSN K0210-1 in NGC 972: 2002 July 26.5 UT, [17.5 (3σ); Nov. 26.3 UT, 15.7; 2003 Jan. 13.1, 16.8 (3σ). W. Li, University of California, infers that examination of KAIT images during 2002 Sept. 5–Dec. 3 shows nothing obvious at the position of K0210-1 (though there may have been a very faint optical transient visible during that time located ∼ 16′′.2 west and 19′′.5 north of the nucleus of NGC 972). Additional approximate K′ magnitudes for PSN K0304-1 in NGC 5900: 2003 May 17.3, 17.9. Li adds that there is no obvious optical transient at the position of K0304-1 in KAIT images taken between 2003 Mar. 5 and July 5 (limiting mag ∼ 21.0).

It is estimated that PSN K0210-1 was first detected no more than 60 days past maximum light and so was never brighter than K′ = 13.4, assuming an average fading of 0.02 mag/day from Mattila and Meikle (2001, MNRAS 342, 325, Table 5); it is anticipated that K0210-1 experienced ∼ 1 mag of extinction in K′. If the limiting mag is 21 near the nucleus in the optical for PSN K0210-1, then these observations suggest ∼ 6 mag of visual extinction or greater toward the infrared object. Assuming that both K0210-1 and K0304-1 have the intrinsic colors of A-type stars, a limiting mag of 21 in the optical suggests at least 3 magnitudes of visual extinction for the source in NGC 5900; for a late-time observation, the source could be intrinsically much redder, requiring less or no extinction. Additional optical observations within a few months of the above observations would be useful to help to set limits on the extinction of each object.

COMET C/2006 OF_{2} (BROUGHTON) Visual total-mag estimates by J. J. Gonzalez, Leon, Spain (20-cm re-

flector): June 28.02 UT, 11.9; Aug. 9.06, 10.8; Sept. 6.98, 9.9; 29.15, 10.0.

2008 October 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green